A. Suppose that a marksman hits the bull's eye 15000 times out of 50000 shots. If the next 4 shots are independent. Find the probability that:
(I)the next 4 hit the bull's eye
(Ii) two of the next shots hits the bull's eye
1
Expert's answer
2020-06-24T18:34:27-0400

The probability that a marksman hits the bull's eye is:

"p=\frac{15000}{50000}=0.3"

(i) We are dealing with 4 independent events. Using the rule of multiplication, we find that the probability is "p^4=(0.3)^4=0.0081"

(ii) Since two successful shots may be at arbitrary places (e.g., first and second shots can be successful, or second and fourth, etc) we have to multiply by the binomial coefficient "C_4^2=\frac{4!}{2!2!}" .(a number of possible subsets of two elements that are chosen from the set of four elements) The resulting probability is "C_4^2p^2(1-p)^2=\frac{4!}{2!2!}(0.3)^2(0.7)^2=0.2646"

Answer: (i) "0.0081" (ii) "0.2646"